# Strategic resource allocation in memory encoding: An efficiency principle shaping language processing

### v1

@@ 2.1.2 Empirical Support
A dynamic similar to strategic resource allocation is observed in the resource-rational model of sentence processing by Hahn et al. (2022). Grounded in the framework of lossy-context surprisal (Futrell, Gibson, and Levy, 2020), their model involves a contextual representation that represents only those words that are most useful for a downstream next-word prediction task. Their model predicts that function words, which are mostly predictable from the linguistic context, are more likely to undergo decay. In fact, our proposal of strategic resource allocation and lossy-context surprisal theory form two sides of the same coin in many aspects. We will discussion the relationship between these two theories in more details in General Discussion.
@@
Studies focusing on the retrieval mechanism in working memory find that linguistic units of higher semantic complexity can be more easily retrieved in later stages of processing despite the initial encoding difficulty, implicating an enhanced accessibility for informative content from the model-theoretic perspective of informativity (Hofmeister, 2011; Hofmeister & Vasishth, 2014; Karimi & Ferreira, 2016; Troyer et al., 2016). However, the exact cognitive underpinning for this empirical observation is still debatable (Hofmeister & Vasishth, 2014; Karimi et al., 2023), and there is lack of clear empirical evidence for whether this effect can be extended to linguistic units of lower predictability (Hofmeister, 2011), a measure that is more consistent with the information-theoretic view of informativity (Shannon, 1948). In spite of these unsettled issues, as a preliminary evidence from the existing literature, the effect of facilitated retrieval for semantically complex units aligns with our rational account outlined above, in the sense that the enhanced accessibility associated with informative units results from the prioritized resources allocated to their encoding.
@@
More recently, strategic resource allocation is more directly examined by Xu & Futrell (2025) through the lens of the agreement attraction effect in English. As shown below, even though the sentences in (1) are ungrammatical in English due to the mismatch of number feature between the subject head noun and the main verb, they are often perceived grammatical by native speakers due to the interference from the distractor noun in between, which shares the number feature with the ungrammatical main verb. In Xu & Futrell (2025), by manipulating the surprisal of the subject head noun through a prenominal adjective, they find that, compared to more surprising subject head nouns (e.g., cute monster), less surprising ones (e.g., evil monster) lead to stronger agreement attraction effect, such that the processing of the main verb is less susceptible to the interference from the distractor noun. They interpret the result as evidence for an enhanced memory representation of more surprising linguistic units against memory interference.
@@
In visual working memory, statistical regularities in long-term knowledge have been shown to shape memory performance. Despite the fact that items more consistent with prior knowledge are easier to be encoded, with lower neural activity and enhanced behavioral performance (Bates and Jacobs, 2020; Blalock, 2015; Jackson & Raymond, 2008; Xie & Zhang, 2017; Girshick et al., 2011), some recent studies indeed observe that, in later stages of processing, these familiar items are de-prioritized to save more resources for the processing of novel ones (Kowialiewski et al., 2022; Bruning and Lewis-Peacock, 2020; Hedayati et al., 2022; Brady et al., 2024). For example, in a delayed-estimation task, Bruning and Lewis-Peacock (2020) ask participants to first memorize and then recall the exact locations of six colored balls on a circle after a brief delay. Before the task, a sub-area on the circle has been previously illustrated to certainly contain the ball with a specific color (e.g., red ball) as a prior information. Their critical finding is that colors not included in the prior information (e.g., non-red balls) have lower recall accuracy when positioned closer to that sub-area, suggesting that memory resources have been shifted away from the prior area to prioritize other areas where novel information is more likely to appear.
@@ 2.2 Dependency Locality
The empirical focus of this paper to exmaine the principle of strategic resource allocation is the locality effect in sentence processing, which has been considered a representative example of the efficient use of working memory resources. Consider the sentence pair in (2). In (2a), codependents in the subject–verb dependency are adjacent to each other, whereas in (2b), there is additional linguistic material in between:
@@
The Dependency Locality Theory (DLT) (Gibson, 1998, 2000) holds that the formation of the non-local structures is constrained by the limited capacity of working memory. Specifically, as dependency distance increases, there is a higher memory cost to store the incomplete dependency as well as a higher integration cost to compute the new structural representation when the later co-dependent is encountered. In support of the DLT, increased processing difficulty is often associated with structures that have longer dependency distance (e.g., Miller & Isard, 1964; Yngve, 1960; Grodner & Gibson, 2005; Bartek et al., 2011; Ford, 1983; King & Just, 1991; Gordon et al., 2001; Traxler et al., 2002). [^1] There is actually an anti-locality effect often found in some head-final dependencies, which is considered to be better explained by an expectation-based mechanism (Vasishth & Lewis, 2006; Nakatani & Gibson, 2010; Konieczny, 2000; Levy & Keller, 2013). Similarly, in the resolution of structural ambiguity where a constituent has multiple potential attachment sites, there is a tendency for comprehenders to prefer the structure with local attachment over others (Frazier and Fodor, 1978; Pearlmutter & Gibson, 2001; Gibson et al., 1996).
@@
These locality principles depend on the precise nature of working memory. Therefore, beyond the general capacity-based constraint proposed by the DLT, it remains an open question how far this efficiency account can go with more and more realistic and detailed characterization of the nature of working memory constraints. Moreover, the existing discussion in the literature rarely addresses efficiency in processing per se. In other words, it is possible that memory limitations make language users not only actively choose a sentence form that is easier to process, but also develop an efficient processing strategy to better handle the information they passively receive.
@@ 2.3 The Current Study
Figure 2: Empirical prediction of strategic resource allocation in dependency locality
We examine strategic resource allocation in the context of dependency locality through naturalistic corpus data. If working memory resources are indeed dynamically and strategically allocated such that novel and unexpected information is prioritized, we predict that antecedents (i.e., left codependents) that are more surprising should receive sharper encoding with less uncertainty. The consequence of this is that memory for more surprising antecedents is enhanced, making their representations less susceptible to memory decay and interference before they need to be re-accessed at the other side of the dependency. Therefore, as illustrated in Figure 2, more surprising antecedents should be able to tolerate longer dependency length, resulting in a reduced locality effect. We approach this prediction from both production (Study 1) and comprehension (Study 2a and 2b).
@@
To preview our results, we find converging evidence from both production and comprehension that unexpected information is encoded with enhanced robustness against decay and interference. In Study 1, which focuses on production data, we observe that more surprising antecedents are associated with longer dependency lengths, an effect that is not reducible to a simple frequency effect. Moreover, the effect mostly exists within Indo-European and head-initial languages in our analysis, and is more consistent for subject relations. In Study 2a and 2b, examining comprehension data from English reading-time corpora, we find a reduced locality effect at the retrieval site for more surprising antecedents. Consistent with Study 1, this effect is more pronounced in subject relations and is observed more reliably in the self-paced reading corpus (Study 2a) than in the eye-tracking corpus (Study 2b).
@@ 3.1.1 Data
    [arc edge, arc angle=85, text only label, label style=above] {deptext}[column sep=1.2em] Isaac & Newton & left & a & note & to & Einstein.
@@
Some UD corpora consist of out-of-context independent sentences, while others are organized document by document, which provide longer and enriched discourse context for each token. This difference may influence our surprisal estimates, which are sensitive to the preceding context of each token. We extracted all the dependencies of a sentences annotated in the UD corpora. All the UD corpora we used, except for Amharic, have a pre-defined split into training, dev, and test sets. We only used the pre-defined training sets since they already have decent sample size.
{adjustwidth}
-0.2in-0.2in Language Corpus Genre # All # Subj # Obj Danish DDT (Johannsen et al., 2015) sent 45,976 4,203 3,963 English GUM (Zeldes, 2017) doc 89,947 7,881 7,296 German GSD (McDonald et al., 2013) sent 155,480 9,602 8,474 Italian ISDT (Bosco et al., 2013) doc 208,939 10,323 11,735 Japanese GSD (Tanaka et al., 2016) sent 113,771 5,005 4,018 Korean Kaist (Chun et al., 2018) doc 154,609 9,855 24,690 Mandarin GSDSimp (Nivre et al., 2020) sent 63,456 5,538 7,576 Russian SynTagRus (Droganova et al., 2018) doc 329,745 32,822 25,065 Spanish AnCora (Taulé et al., 2008) doc 333,728 21,472 31,143 Turkish BOUN (Marşan et al., 2022) sent 45,914 3,861 4,680
@@ 3.1.2 Estimating Token Surprisal
In this work, to ensure that the results are not the artifact of a specific language model, we generated surprisal measures from both the GPT-3 base (text-davinci-001; Brown et al., 2020) and the mGPT language models (Shliazhko et al., 2024), both being trained on multilingual data. For each token $w$ in the dependency corpora, we obtained its surprisal $-\log p(w_{t}\mid w_{<t})$ given the preceding context from both models. We used the maximally allowed context window in the corresponding document or sentence. It is worth noting that Mandarin Chinese, unfortunately, is not supported by mGPT. Therefore, we only report the results with GPT-3 surprisal for Mandarin.
@@ 3.1.3 Measuring Dependency Length
We did the analysis with two different measures of dependency length $L$. The first measure is an orthographic one $L_{\mathrm{O}}$, which is the number of words between the codependents of a dependency. The second measure is an information-theoretic one $L_{\mathrm{I}}$, which sums up the surprisal of all the words from $w_{i}$ to $w_{i+N}$:
@@ 3.1.4 Data Transformation and Exclusion
Constructions such as foreign phrases, multi-word proper names, and fixed expressions are annotated as flat structures in UD corpora. We merged flat structures such that the surprisal of the whole structure is the sum of all its components, and that the first word in the flat structure is treated as the head when calculating the length of a dependency involving a flat structure. For example, the subject–verb dependency in (4) involves a flat structure in the subject position. The antecedent surprisal for this dependency is thus the sum of surprisal over both words “Isaac Newton,” and the dependency length by word counts is one since the first word “Isaac” is one word away from the verb. We excluded sentences that are less than five-word long, since sentences that are too short may have limited room for the dependency length to vary and many of the short “sentences” are in fact titles and extended proper names (e.g., e-mail addresses and institution names). We excluded punctuation tokens. We also excluded tokens whose surprisal value is greater than 20 bits, as the surprisal estimates for such rare word sequences may be unreliable. Moreover, exceedingly surprising information may introduce confounding factors in human processing. We then extracted all the dependencies in which both the head and the dependent are spared from data exclusion.
@@ 3.1.5 Data Analysis
For the analysis with the full dataset, for each language, we ran separate linear mixed-effects models predicting the two variants of dependency length $L$ as the dependent variable, using the lmerTest package in R (Kuznetsova et al., 2017). The critical fixed-effect predictor is the antecedent surprisal, with random intercept by dependency types. [^7] As mentioned above, we also compare the effect of antecedent surprisal with antecedent frequency in the current analysis. However, the models with random slopes for both effects rarely converge. Therefore, for better interpretability of the statistical result, we only included random intercept by dependency types. For the analyses with subject and object relations, we ran linear models with the same fixed effects. We included four control variables for all the analyses, as in (5). Sentence position aims to control the discourse-level information structure, where more information may be given as the discourse develops; Antecedent position aims to control that antecedents appearing towards to the end of a sentence naturally tend to have shorter dependency length. Sentence length aims to control for two possible confounds: first, longer sentences may tend to have longer dependency length in general; second, longer sentences may tend to have more complex syntactic structure, which may be associated with more surprising antecedents. We also included antecedent frequency in log scale retrieved from Speer (2022) in order to see whether the surprisal effect is reducible to a simple frequency effect as mentioned above. All variables are $z$-scaled.
@@
    • sentence position: position of the sentence in the current document, if the corpus is organized document-by-document
@@ 3.2 Result
Figure 4: Study 1 coefficient estimates for the effects of antecedent surprisal and antecedent frequency on dependency length $L$ across languages, with 95% confidence interval. Our hypothesis of strategic resource allocation predicts that more surprising antecedents are associated with longer $L$ (positive effect of antecedent surprisal), and that more frequent antecedents are associated with shorter $L$ (negative effect of antecedent frequency). Significance levels: *($p$$<$0.05), **($p$$<$0.01), ***($p$$<$0.001)
The result of the raw data in its original scale is presented in Figure 3, which shows dependency length $L$ as a function of antecedent surprisal. The statistical result of the regression models is presented in Figure 4 for the effects of antecedent surprisal and antecedent frequency. A high-level summary of the statistical evidence across languages is shown in Figure 5 for the critical effects. Since we used the surprisal measure from GPT-3 and mGPT, we describe an effect as robust and independent of model parameterization if it is significant in the same direction both in the analysis with GPT-3 and in the one with mGPT (i.e., both positive or both negative, highlighted in dark red and dark blue in Figure 5). [^8] Since Mandarin is not available for mGPT, a critical effect is highlighted in dark red or blue in Figure 5 even though we only have the result with GPT-3. We describe the effect as partially confirmed and less robust if it reaches significance with only one of the language models (highlighted in light red and blue in Figure 5). We describe the effect as inconclusive if it is not significant with any model, or if GPT-3 and mGPT show significantly conflicting result (i.e., significant positive in one model but significant negative in the other). [^9] The use of these terms (i.e., robust, less robust, partially confirmed, and inconclusive) is only for expository purpose, and does not imply any direct statistical test.
Figure 5: Study 1 summary of statistical result for the effects of antecedent surprisal and antecedent frequency on dependency length $L$. Significant ($p$$<$0.05) positive effects are highlighted in red; significant negative effects are highlighted in blue. Effects are considered not conclusive if insignificant with both language models, or if GPT-3 and mGPT show conflicting result where the effect is significant in opposite directions. According to our hypothesis, antecedent surprisal is expected to have a positive effect on $L$, whereas antecedent frequency is expected to show negative effect.

### v2

@@ 2.4 Some Existing Empirical Evidence
A dynamic similar to strategic resource allocation (SRA) is observed in the resource-rational model of sentence processing by Hahn et al. (2022). Grounded in the framework of lossy-context surprisal (Futrell, Gibson, and Levy, 2020), their model involves a contextual representation that represents only those words that are most useful for a downstream next-word prediction task. Their model predicts that function words, which are mostly predictable from the linguistic context, are more likely to undergo decay. In fact, our proposal of strategic resource allocation and lossy-context surprisal theory form two sides of the same coin in many aspects. We will discuss the relationship between these two theories in General Discussion Section 7.4.
@@
Studies focusing on memory retrieval mechanisms find that linguistic units of higher semantic complexity can be more easily retrieved in later stages of processing despite the initial encoding difficulty, implicating an enhanced accessibility for informative content from the model-theoretic perspective of informativity (Hofmeister, 2011; Hofmeister & Vasishth, 2014; Karimi & Ferreira, 2016; Troyer et al., 2016). However, the exact cognitive underpinning for this empirical observation is still debatable (Hofmeister & Vasishth, 2014; Karimi et al., 2023), and there is lack of clear empirical evidence for whether this effect can be extended to the information-theoretic view of informativity based on probabilistic prediction (Shannon, 1948). In spite of these unsettled issues, as a preliminary evidence from the existing literature, the effect of facilitated retrieval for semantically complex units aligns with our rational account outlined above, in the sense that the enhanced accessibility associated with informative units results from the prioritized resources allocated to their encoding.
@@
Recently, SRA is more directly examined by Xu & Futrell (2025) through the lens of the agreement attraction effect in English. As shown below, even though the sentences in (3) are ungrammatical in English due to the mismatch of number feature between the subject head noun and the main verb, they are often perceived grammatical by native speakers due to the interference from the distractor noun in between, which shares the number feature with the ungrammatical main verb. In Xu & Futrell (2025), by manipulating the surprisal of the subject head noun through a prenominal adjective, they find that, compared to more surprising subject head nouns (e.g., cute monster), less surprising ones (e.g., evil monster) lead to stronger agreement attraction effect, such that the processing of the main verb is less susceptible to the interference from the distractor noun. They interpret the result as evidence for an enhanced memory representation of more surprising linguistic units against memory interference.
@@
In visual working memory, statistical regularities in long-term knowledge have been shown to shape memory performance. Despite the fact that items more consistent with prior knowledge are easier to be encoded with lower neural activity and enhanced behavioral performance (Bates and Jacobs, 2020; Blalock, 2015; Jackson & Raymond, 2008; Xie & Zhang, 2017; Girshick et al., 2011), some recent studies indeed observe that, in later stages of processing, these familiar items are de-prioritized to save more resources for the processing of novel ones (Kowialiewski et al., 2022; Bruning and Lewis-Peacock, 2020; Hedayati et al., 2022; Brady et al., 2024). For example, in a delayed-estimation task, Bruning and Lewis-Peacock (2020) ask participants to first memorize and then recall the exact locations of six colored balls on a circle after a brief delay. Before the task, a sub-area on the circle has been previously illustrated to certainly contain the ball with a specific color (e.g., red ball) as a prior information. Their critical finding is that colors not included in the prior information (e.g., non-red balls) have lower recall accuracy when positioned closer to that sub-area, suggesting that memory resources have been shifted away from the prior area to prioritize other areas where novel information is more likely to appear.
3 SRA and Dependency Locality
The empirical focus of this paper to examine SRA is the locality effect in sentence processing, which has been considered a representative example of the efficient use of working memory resources. In this section, we will first introduce the empirical background of dependency locality effect. Then, we will present the empirical predictions of SRA in the context of dependency locality.
@@ 3.1 Dependency Locality
Consider the sentence pair in (4). In (4a), codependents in the subject–verb dependency are adjacent to each other, whereas in (4b), there is additional linguistic material in between:
@@
The Dependency Locality Theory (DLT) (Gibson, 1998, 2000) holds that the formation of the non-local structures is constrained by the limited capacity of working memory. Specifically, as dependency distance increases, there is a higher memory cost to store the incomplete dependency as well as a higher integration cost to compute the new structural representation when the other codependent is encountered. In support of DLT, increased processing difficulty is often associated with structures that have longer dependency distance (e.g., Miller & Isard, 1964; Yngve, 1960; Grodner & Gibson, 2005; Bartek et al., 2011; Ford, 1983; King & Just, 1991; Gordon et al., 2001; Traxler et al., 2002). [^7] There is actually an anti-locality effect often found in some head-final dependencies, which is considered to be better explained by an expectation-based mechanism (Vasishth & Lewis, 2006; Nakatani & Gibson, 2010; Konieczny, 2000; Levy & Keller, 2013). Similarly, in the resolution of structural ambiguity where a constituent has multiple potential attachment sites, there is a tendency for comprehenders to prefer the structure with local attachment (Frazier and Fodor, 1978; Pearlmutter & Gibson, 2001; Gibson et al., 1996).
@@
These locality principles depend on the precise nature of working memory. Therefore, beyond the general capacity-based constraint proposed by DLT, it remains an open question how far this efficiency account can go with more and more realistic and detailed characterization of the nature of working memory constraints. Moreover, the existing discussion in the literature rarely addresses efficiency in processing per se. In other words, it is possible that memory limitations make language users not only actively choose a sentence form that is easier to process, but also develop an efficient processing strategy to better handle the information they passively receive.
@@ 3.2 The Current Study
Figure 4: Empirical prediction of strategic resource allocation in dependency locality. High-surprisal antecedents are more tolerable to longer dependency length.
We examine SRA in the context of dependency locality through naturalistic corpus data. If working memory resources are indeed dynamically and strategically allocated such that novel and unexpected information is prioritized, we predict that antecedents (i.e., left codependents) that are more surprising should receive sharper encoding with less uncertainty. The consequence of this is that memory for more surprising antecedents is enhanced, making their representations less susceptible to memory decay and interference before they need to be re-accessed at the other side of the dependency. Therefore, as illustrated in Figure 4, more surprising antecedents should be able to tolerate longer dependency length, resulting in a reduced locality effect. We approach this prediction from both production (Study 1) and comprehension (Study 2a and 2b).
@@
To preview our results, we find converging evidence from both production and comprehension that unexpected information is encoded with enhanced robustness against decay and interference. In Study 1, which focuses on production data, we observe that more surprising antecedents are associated with longer dependency lengths, an effect that is not reducible to a simple frequency effect. Moreover, the effect mostly exists within Indo-European and head-initial languages in our analysis, and is more consistent for subject relations. We discuss the cross-linguistic variability in General Discussion. In Study 2a and 2b, examining comprehension data from English reading-time corpora, we find a reduced locality effect at the retrieval site for more surprising antecedents. Consistent with Study 1, this effect is more pronounced in subject relations and is observed more reliably in the self-paced reading corpus (Study 2a) than in the eye-tracking corpus (Study 2b).
@@ 4.1.1 Data
    [arc edge, arc angle=85, text only label, label style=above] {deptext}[column sep=0.8em] Isaac & Newton & left & a & note & to & Einstein.
@@
Some UD corpora consist of out-of-context independent sentences, while others are organized document by document, which provide longer and enriched discourse context for each token. This difference may influence our surprisal estimates, which are sensitive to the preceding context of each token. We extracted all the dependencies of a sentence annotated in the UD corpora. All the UD corpora we used have a pre-defined split into training, dev, and test sets. We only used the pre-defined training sets since they already have decent sample size.
Language Corpus Genre # All # Subj # Obj
Danish DDT (Johannsen et al., 2015) sent 45,976 4,203 3,963
English GUM (Zeldes, 2017) doc 89,947 7,881 7,296
German GSD (McDonald et al., 2013) sent 155,480 9,602 8,474
Italian ISDT (Bosco et al., 2013) doc 208,939 10,323 11,735
Japanese GSD (Tanaka et al., 2016) sent 113,771 5,005 4,018
Korean Kaist (Chun et al., 2018) doc 154,609 9,855 24,690
Mandarin GSDSimp (Nivre et al., 2020) sent 63,456 5,538 7,576
Russian SynTagRus (Droganova et al., 2018) doc 329,745 32,822 25,065
Spanish AnCora (Taulé et al., 2008) doc 333,728 21,472 31,143
Turkish BOUN (Marşan et al., 2022) sent 45,914 3,861 4,680
@@ 4.1.2 Estimating Token Surprisal
In this work, to ensure that the results are not the artifact of a specific language model, we generated surprisal measures from both the GPT-3 base (text-davinci-001; Brown et al., 2020) and the mGPT language models (Shliazhko et al., 2024), both being trained on multilingual data. For each token $w$ in the dependency corpora, we obtained its surprisal $\log p(w_{t}\mid w_{<t})$ given the preceding context from both models. We used the maximally allowed context window in the corresponding document or sentence. It is worth noting that Mandarin Chinese, unfortunately, is not supported by mGPT. Therefore, we only report the results with GPT-3 surprisal for Mandarin.
@@ 4.1.3 Measuring Dependency Length
We did the analysis with two different measures of dependency length $L$. The first measure is an orthographic one $L_{\mathrm{O}}$, which is the number of words between the codependents of a dependency. The second measure is an information-theoretic one $L_{\mathrm{I}}$, which sums up the surprisal of all words between codependents from $w_{i}$ to $w_{i+N}$:
@@ 4.1.4 Data Transformation and Exclusion
Constructions such as foreign phrases, multi-word proper names, and fixed expressions are annotated as flat structures in UD corpora. We merged flat structures such that the surprisal of the whole structure is the sum of all its components, and that the first word in the flat structure is treated as the head when calculating the length of a dependency. For example, the subject–verb dependency in (6) involves a flat structure in the subject position. The antecedent surprisal for this dependency is thus the sum of surprisal over both words “Isaac Newton,” and the dependency length by word counts is 1, since the first word “Isaac” is one word away from the verb. We excluded sentences that are less than five-word long, since sentences that are too short may have limited room for the dependency length to vary and many of the short “sentences” are in fact titles and extended proper names (e.g., e-mail addresses and institution names). We excluded punctuation tokens. We also excluded tokens whose surprisal value is greater than 20 bits, as the surprisal estimates for such rare word sequences may be unreliable. Moreover, exceedingly surprising information may introduce confounding factors in human processing. We then extracted all the dependencies in which both the head and the dependent are spared from data exclusion.
@@ 4.1.5 Data Analysis
For the analysis with the full dataset, for each language, we ran separate linear mixed-effects models predicting the two variants of dependency length $L$ as the dependent variable, using the lmerTest package in R (Kuznetsova et al., 2017). The critical fixed-effect predictor is the antecedent surprisal, with random intercept by dependency types. [^13] As mentioned above, we also compare the effect of antecedent surprisal with antecedent frequency in the current analysis. However, the models with random slopes for both effects rarely converge. Therefore, for better interpretability of the statistical result, we only included random intercept by dependency types. For the analyses with subject and object relations, we ran linear models with the same fixed effects. We included five control variables for all the analyses, as in (7). Sentence position aims to control the discourse-level information structure, where more information may be given as the discourse develops; Antecedent position aims to control that antecedents appearing towards to the end of a sentence naturally tend to have shorter dependency length. Sentence length aims to control for two possible confounds: first, longer sentences may tend to have longer dependency length in general; second, longer sentences may tend to have more complex syntactic structure, which may be associated with more surprising antecedents. We also included antecedent frequency in log scale retrieved from Speer (2022) in order to see whether the surprisal effect is reducible to a simple frequency effect. For the analysis with information-theoretic dependency length $L_{\mathrm{I}}$, we included an additional control variable baseline surprisal, which is the surprisal averaged across all words within a sentence. This is to address the confound that sentences with higher baseline surprisal naturally leads to a positive correlation between antecedent surprisal and the information-theoretic $L_{\mathrm{I}}$. All variables are $z$-scaled.
@@
    • sentence position: position of the sentence in the current document (only included if the corpus is organized document-by-document)
@@
    • baseline surprisal: average surprisal across all words within a sentence (only included for the analysis with information-theoretic dependency length $L_{\mathrm{I}}$)
@@ 4.2 Result
Figure 6: Study 1 coefficient estimates for the effects of antecedent surprisal and antecedent frequency on dependency length $L$ across languages, with 95% confidence interval. Our hypothesis of strategic resource allocation predicts that more surprising antecedents are associated with longer $L$ (positive effect of antecedent surprisal), and that more frequent antecedents are associated with shorter $L$ (negative effect of antecedent frequency). Significance levels: *(pp<<0.05), **(pp<<0.01), ***(pp<<0.001)
The result of the raw data in its original scale is presented in Figure 5, which shows dependency length $L$ as a function of antecedent surprisal. The statistical result of the regression models is presented in Figure 6 for the effects of antecedent surprisal and antecedent frequency. A high-level summary of the statistical evidence across languages is shown in Figure 7. Since we used the surprisal measure from GPT-3 and mGPT, we describe an effect as robust and independent of model parameterization if it is significant in the same direction both in the analysis with GPT-3 and in the one with mGPT (i.e., both positive or both negative, highlighted in dark red and dark blue in Figure 7). [^14] Since Mandarin is not available for mGPT, a critical effect is highlighted in dark red or blue in Figure 7 even though we only have the result with GPT-3. We describe the effect as partially confirmed and less robust if it reaches significance with only one of the language models (highlighted in light red and blue in Figure 7). We describe the effect as inconclusive if it is not significant with any model, or if GPT-3 and mGPT show significantly conflicting result (i.e., significantlt positive in one model but significantly negative in the other). [^15] The use of these terms (i.e., robust, less robust, partially confirmed, and inconclusive) is only for expository purpose, and does not imply any direct statistical robustness test.
Figure 7: Study 1 summary of statistical result for the effects of antecedent surprisal and antecedent frequency on dependency length $L$. Significant (pp<<0.05) positive effects are highlighted in red; significant negative effects are highlighted in blue. Effects are considered not conclusive if insignificant with both language models, or if GPT-3 and mGPT show conflicting result where the effect is significant in opposite directions. According to our hypothesis, antecedent surprisal is expected to have a positive effect on $L$, whereas antecedent frequency is expected to show negative effect.